# Laminar chaos in systems with random and chaotically time-varying delay

### v1

@@ IV.1 Numerical results
Figure 3: Influence of the random delays shown in Fig. 1 on the dimension of chaotic attractors: In the top panel a heat map of Lyapunov exponent $\lambda[R]$ of the access map, Eq. (10), is shown as a function of the mean delay $\tau_{0}$ and the delay amplitude $A$. For fixed $A=0.9$ (white dashed line), $\lambda[R]$ and the Kaplan-Yorke dimension $D_{KY}$ of Eq. (1) is shown in the center and bottom panels (blue solid lines), respectively. While for smaller $\tau_{0}$ we find fractal structures similar to those in Fig. 4, for increasing $\tau_{0}$ the parameter space becomes more and more regular, where only dissipative delays, $\lambda[R]<0$ appear. Therefore, in almost the whole parameter space, a drastic reduction of $D_{KY}$ compared to a system with constant delay (black solid line, see text) is observed. Laminar chaos is found for $\lambda[f]+\lambda[R]<0$, i.e., whenever $\lambda[R]$ is below the dotted line in the center panels.
@@
In the following, we consider Eq. (1) with a quadratic nonlinearity $f(z)=3.8\,z(1-z)$ and the random and chaotically time-varying delays illustrated in Fig. 1 and Fig. 2(b), respectively. For the numerical analysis we choose $\Theta=200$ such that the small time scale $(1/\Theta)$ of the system is much smaller than the delay $\tau(t)$ and also much smaller than the time scale of the temporal structures of the delay variation. The results are presented in Figs. 3 and 5 for random and chaotically time-varying delays, respectively. For highlighting the differences between these and regular delay variations, analogous results for a periodic (sinusoidal) delay variation are shown in Fig. 4. Since laminar chaos and its generalization can only be observed in systems with dissipative delay, our first step is the classification of the time-varying delays in the delay parameter space using the Lyapunov exponent $\lambda[R]$ of the access map. For general one-dimensional maps, the Lyapunov exponent is defined by [105]
@@
To highlight the differences that come with random and chaotically time-varying delay, we first consider a generic example of a regular delay variation: The sinusoidal delay given by Eq. (2). Its parameter space shown in the top panel of Fig. 4 is characterized by two dynamical regimes of the access map and its corresponding circle map, Eq. (16) with $L=1$. Inside the so-called Arnold tongues [112, *arnold_small_1961_erratum], which are the dark tongue-like structures where $\lambda[R]$ is negative, the circle map shows stable periodic dynamics and the resulting delays are classified as dissipative delays. The complement of the Arnold tongues, which for fixed $A$ is a Cantor set with nonzero measure, i.e., a fat fractal [105], leads to marginally stable quasiperiodic dynamics of the circle map, $\lambda[R]=0$ and thus the resulting delay are classified as conservative delays. As illustrated in the center panel, $\lambda[R]$ as a function of $\tau_{0}$ for fixed $A=0.9$ also reflects this fractal structure showing an irregular behavior while it is periodic in $\tau_{0}$, where the period equals the delay period. For the delay system this means that the dynamics changes irregularly between high-dimensional turbulent chaos, and low-dimensional (generalized) laminar chaos if the delay parameters are varied. This has drastic consequences for the Kaplan-Yorke dimension $D_{KY}$ as shown in the bottom panel for fixed $A=0.9$. If the mean delay $\tau_{0}$ is varied, one observes huge variations of $D_{KY}$ over several orders of magnitudes, where the local maxima are roughly proportional to $\tau_{0}$. The overall variation of $D_{KY}$ increases with the parameter $\Theta$ since the local maxima are also proportional to $\Theta$ as they correspond to conservative delays leading to turbulent chaos, where the system is equivalent to a constant delay system so that the theory in [74] applies. In contrast, the local minima hardly vary with $\Theta$ given that $\Theta$ is large enough since the system is then well described by the low-dimensional dynamics of the limit map, Eq. (12). Such huge variations of $D_{KY}$ in a fractal manner are also observed for quasiperiodic delays [54].
@@
The situation drastically changes, when it comes to a random delay variation as shown in Fig. 3. While for small values of the mean delay $\tau_{0}$ we still find fractal structures in the parameter space similar to the ones observed for periodic delays, the fractal structures disappear for larger $\tau_{0}$. Moreover, the Lyapunov exponent of the access map seems strictly negative for all $A>0$ if $\tau_{0}$ exceeds a certain threshold, which indicates that almost all parameters lead to dissipative delays. This is also confirmed in the center panels, where $\lambda[R]$ is shown as a function of $\tau_{0}$ for fixed $A=0.9$. According to the theory on laminar chaos in Sec. III, the delay system defined by Eq. (1) shows low-dimensional (generalized) laminar chaos for almost all delay parameters, when $\tau_{0}$ exceeds a certain threshold. In Sec. IV.2 we find that this threshold is of the order of the correlation length of the delay variation. The low-dimensionality of the dynamics of the delay system is confirmed by the bottom panels, where we find that the Kaplan-Yorke dimension $D_{KY}$ for random delays (blue line) with large enough $\tau_{0}$ is of orders of magnitudes smaller than the dimension observed for the comparison system with a constant delay (black line).
@@ IV.2 Theory
In order to explain the strictly negative value of the Lyapunov exponent of the access map that is observed in Fig. 3 and Fig. 5 for larger values of $\tau_{0}$, we follow the argumentation on disordered circle maps in the limit $L\to\infty$ found in [94]. There the finite-time Lyapunov exponent $\lambda_{K_{L}}[R_{\omega}]=\frac{1}{K_{L}}\sum_{k=0}^{K_{L}-1}\ln(1-\dot{\tau}_{\omega}(\theta_{k}))$ of an orbit $\mathcal{O}_{L}=\{\theta_{k}\}_{k}=0,1,\dots,K_{L}-1$ that passes the phase space of the circle map once is considered. Assuming that the minimal delay $\tau_{\text{min}}=\min_{\omega,t}\tau_{\omega}(t)$ is larger than the correlation length $l$ of the delay $\tau_{\omega}(t)$, $\tau_{\text{min}}>l$, one finds for the delays defined in Sec. II that the increments $\tau_{\omega}(\theta_{k})$ and $\tau_{\omega}(\theta_{k+1})$ are nearly independent. More precisely, if the tails of the the Gaussian $G_{\varsigma}(t)$ in Eq. (7) are neglected and therefore the Gaussian is replaced by a distribution with finite support $[-l,l]$ the increments are strictly independent. In this case, the circle map with quenched disorder defined by Eq. (20) is stochastically equivalent to the circle map with annealed disorder, where the disorder changes with each iteration $\omega=\omega_{k}$. In the literature, such maps are known as “random dynamical system on the circle” [117, 118] (see [94] for more references) and it is known that their Lyapunov exponent is strictly negative if the realizations of the circle maps obtained from the disorder realizations $\omega_{k}$, almost surely, do not preserve the same measure [117, 118], which is a very general assumption. This means that $\lambda_{K_{L}}[R_{\omega}]$ converges to a unique, strictly negative value $\lambda[R]$ in the limit $L\to\infty$, which is associated with $K_{L}\to\infty$ as the length $K_{L}$ of the orbit $\mathcal{O}_{L}$ grows with the system size, $K_{L}\sim L/T$, where $T$ is the limit of the average roundtrip time, Eq. (22), for $L\to\infty$. For finite $L$, the authors of [94] further consider the distribution of the Lyapunov exponent of the disordered circle map, Eq. (20), with respect to the disorder. Using the approximate independence of subsequent increments $\tau_{\omega}(\theta_{k})$ and the central limit theorem, they find that the probability of generating a disorder realization with a Lyapunov exponent larger than $\lambda_{\text{cutoff}}$ with $\lambda_{\text{cutoff}}>\lambda[R]$ vanishes for increasing $L$ with
@@ V.1 Test for laminar chaos
Figure 7: Return maps of the levels of the laminar phases. The level $z_{n+p}$ of the $(n+p)$th laminar phase is plotted over the level $z_{n}$ of the $n$th laminar phase for different $p$. The $z_{n}$ were extracted from a laminar chaotic time series of Eq. (1) with the chaotically time varying delay shown in Fig. 2(b), where we set $\tau_{0}=3$ and $A=0.9$. While non of the return maps resembles the nonlinearity of Eq. (1) (dashed lines), remnants of the nonlinearity are visible for $1\leq p\leq 4$, which indicates that $p$ is time-dependent and varies between these values for the given parameters.
Figure 8: Analysis of the nonlinear correlations in the return maps in Fig. 7 of a laminar chaotic time series. A value of ${(1-Q)}$ and the MIC close to one would indicate strong nonlinear correlations as one would expect for at least one value of $p$ in the case of laminar chaos in systems with (quasi-)periodic delay. For random and chaotically time-varying delays, much smaller values are observed for all $p$, which indicates only weak correlations. This supports our conjecture that $p$ is time-dependent in this case.
@@
In the following we naively apply the original toolbox to a time series of a system with chaotically time-varying delay, where we choose the same parameters as in Fig. 5 and Fig. 6(d)  [^5] The time series has a total length of $10^{4}$ time units, where we skip the first $10^{3}$ time units to let the transients relax. . We first identify the laminar phases and extract their levels $z_{n}$  [^6] To identify the laminar phases, we first detect the bursts between the laminar phases, where a burst is assumed when the absolute valued derivative $|\dot{z}(t)|$ of the time series exceeds the threshold $0.1$ and the duration between the end of the preceding burst and the current burst candidate is larger than $0.05$. So a laminar phase is detected if $|\dot{z}(t)|$ is close to zero for a duration that is at least one magnitude larger than the characteristic time scale $1/\Theta$ with $\Theta=200$. The burst position is then given by the onset of the burst, where $|\dot{z}(t)|$ starts to exceed the threshold. The level $z_{n}$ of the $n$th laminar phase is defined as the value of $z(t)$ in the middle between two subsequent burst positions.. Then the return map of the laminar phases is plotted, which should reveal the nonlinearity $f$ of the delay system if laminar chaos is present. However, this concept is only partially applicable for chaotically time-varying delay as shown in Fig. 7. The return maps clearly show remnants of the nonlinearity $f$ but with a noisy background and with remnants of higher iterates $f^{l}$. While for periodic and quasiperiodic delay the first iteration of the map $z^{\prime}=f(z)$ appears for a unique value $p=p_{\text{true}}$, it is found in Fig. 7 for $p=1,2,3,4$ (dashed line). This indicates that the number $p_{\text{true}}$ of laminar phases per solution segment is in general not constant for random and chaotically time-varying delay. As done in [95], we now quantify the nonlinear correlations of these return maps, where strong nonlinear correlations are an indicator for the presence of laminar chaos generated from systems with periodic or quasiperiodic delay. For that we compute the maximal information coefficient (MIC) introduced in [127] using minepy [128] and the quantity $(1-Q)$, which relates local fluctuations of the return map to global fluctuations of the levels $z_{n}$  [^7] Following the idea of the CANOVA method [141], we relate local fluctuations of the return maps of the $z_{n}$ to their global fluctuations. Therefore we sort the sequence of points $(z_{n},z_{n+p})$ by the first component leading to the sequence of points $(v_{m},w_{m})$. The quantity $Q$ is then defined by $Q=\frac{\text{Var}(w_{m+1}-w_{m})}{2\,\text{Var}(w_{m})}$, where $\text{Var}(w_{m})$ is the sample variance of all values $w_{m}$. More details can be found in [95]. . The results for our system with chaotic delay are shown in Fig. 8, where values close to 00 or $1$ belong to weak or strong correlations, respectively. Since both quantities are rather close to 00 than to $1$, the classical test for laminar chaos fails. The reason for that is illustrated in Fig. 9, where part of the analyzed time series is shown together with the numerically detected boundaries of the laminar phases and with the boundaries of the state intervals $\mathcal{I}_{k}=(t_{k-1},t_{k}]=(t_{k}-\tau(t_{k}),t_{k}]$, which are the domains of the solution segments $z_{k}(t)$ representing the memory of the delay system at time $t=t_{k}$. Highlighted by the different shadings, the number $p=p_{k}$ of laminar phases per solution segment varies, so that the original test for laminar chaos fails as $p$ is assumed constant. If the delay function $\tau(t)$ is known, it is straight forward to obtain the sequence $p_{n}$ such that the points $(z_{n},z_{n+p_{n}})$ resemble the nonlinearity $f$. Therefore we assign to the $n$th laminar phase a time $\tilde{t}_{n}$ equal to the numerically detected center of the phase. Using the monotonicity of the access map, we know that the $n$th laminar phase is uniquely mapped to the $(n+p_{n})$th laminar phase. It follows that the correct $p_{n}$ with $z_{n+p_{n}}=f(z_{n})$ is found if $R(\tilde{t}_{n+p_{n}})$ is inside the domain of the $n$th laminar phase. Applying this approach to our exemplary time series gives us the associated sequence $p_{n}$. However, for some $n$ no or no unique $p_{n}$ can be detected. This happens when the plateau detection algorithm fails to detect a burst between two laminar phases, when spurious laminar phases are detected, or naturally when the number of laminar phases per solution segment changes, $p_{n+1}\neq p_{n}$ (see transitions between the regions with different color in Fig. 9). Plotting the points $(z_{n},z_{n+p_{n}})$ while omitting points with missing $p_{n}$ results in the return map shown in Fig. 10, clearly reproducing the nonlinearity except for some outliers and showing strong nonlinear correlations as expected for laminar chaos.
Figure 11: Mechanism behind laminar chaos leads to characteristic structures in $d$-dimensional return maps of the durations of the laminar phases with (a) $d=2$, (b) $d=3$. Tuples of durations $(\delta_{n},\delta_{n+1},\dots)$ of laminar phases with the same number $p_{n}=p_{n+1}=\dots=p$ of laminar phases inside the solution segments belong to the same point cluster. Depending on $p$ these clusters scatter around $(p-1)$-dimensional affine subspaces, which can be used to reconstruct the sequence $p_{n}$ from time series without knowing the system (see text). In (a), $p=1$, $p=2$, $p=3$, $p>3$ are represented by the blue squares, orange triangles, green diamonds, and light gray dots, respectively. If no unique $p$ is numerically found, dark grey crosses are used. In (b), the color coding is same and the $(p-1)$ subspaces defining the shape of the cluster are illustrated for $p=1,2,3$ by the black dot, line, and plane, respectively.
Figure 12: Quantity $(1-Q)$ as a function of the parameter $T$ of the clustering algorithm. For each $T$, the sequence $p_{n}$ is extracted from the $(d=4)$-dimensional durations return map. Then $(1-Q)$ is computed for the points $(x_{n},x_{n+p_{n}})$. The maximum at $T=2.78$ (dashed line) is located within the range of the time-varying delay (dotted lines).
@@
Figure 14: Return map of the levels $z_{n}$ of the laminar phases, where the sequence of the $p_{n}$ was extracted from the time series without knowing the system but using the clusters identified in Fig. 13. The nonlinearity (dashed line) is well resembled, but we find more outliers compared to Fig. 10, where the delay function was used to determine the sequence of the $p_{n}$. This is also reflected by slightly lower values of the quantities $(1-Q)\approx 0.79$ and $\text{MIC}\approx 0.77$, which nevertheless indicate strong nonlinear correlations as expected for laminar chaos.
If the system and therefore the time varying delay $\tau(t)$ is unknown, the sequence $p_{n}$ in principle can be reconstructed from the durations $\delta_{n}$ of the laminar phases using the $d$-dimensional return maps given by the points $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$. In Fig. 11, the two and three-dimensional return maps of the durations $\delta_{n}$ of the laminar phases are shown, which were computed by subtracting subsequent burst positions obtained from our exemplary time series. The color coding of the values of the $p_{n}$ from our preceding analysis reveals characteristic structures that can be exploited for obtaining the $p_{n}$ even when we do not know anything about the system. The structures arise from this trivial fact: Given that a solution segment $z_{k}(t)$ contains only complete laminar phases, then the sum of the durations of these laminar phases are equal to the length of its domain, which is the state interval $\mathcal{I}_{k}=(t_{k}-\tau(t_{k}),t_{k}]$ and is of one delay length $\tau(t_{k})$. If there are $p$ laminar phases inside the solution segments, it follows that the components of each point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ of our $d$-dimensional return map fulfill $(d-p+1)$ equations given by
$T_{l}=\sum_{m=l}^{l+(p-1)}\delta_{n+m},\quad l=0,1,\dots,d-p.$ (25)
So each $T_{l}$ is simply the sum of $p$ subsequent durations starting with $\delta_{n+l}$. Considering $T_{l}$ as constants and the durations $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ as variables, Eqs. (25) define a $(p-1)$-dimensional affine subspace of $\mathbb{R}^{d}$, where the $T_{l}$ define its offset from the origin. Since we assumed that $p$ is equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$, each $T_{l}$ is a length of a state interval and therefore is equal to a value of the time-varying delay $\tau(t)$. This means that each vector composed of $d$ subsequent durations belongs to an $(p-1)$-dimensional affine subspace, where the offset of the subspace fluctuates according to the time-varying delay $\tau(t)$. Averaging these fluctuations by replacing $T_{l}$ with the average roundtrip time of the feedback loop $T=T_{\omega}$ from Eq. (22) with $L\to\infty$ leads to the subspaces represented by the black dot, line, and plane in Fig. 11(b), where the clusters with the corresponding values of $p=1,2,3$ scatter around. For our chaotically time-varying delay with the parameters used here, we obtained an average roundtrip time of $T=T_{\omega}=3.0.$ The loose points away from these clusters belong to the case where $p$ is not equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$, which happens at a transition from one value of $p_{n}$ to another, see Fig. 9.

### v2

@@ IV.1 Numerical results
Figure 3: Influence of the random delays shown in Fig. 1 on the dimension of chaotic attractors: In the top panel a heat map of Lyapunov exponent $\lambda[R]$ of the access map, Eq. (10), is shown as a function of the mean delay $\tau_{0}$ and the delay amplitude AA. For fixed $A=0.9$ (white dashed line), $\lambda[R]$ and the Kaplan-Yorke dimension $D_{KY}$ of Eq. (1) is shown in the center and bottom panels (blue solid lines), respectively. While for smaller $\tau_{0}$ we find fractal structures similar to those in Fig. 4, for increasing $\tau_{0}$ the parameter space becomes more and more regular, where only dissipative delays, $\lambda[R]<0$ appear. Therefore, in almost the whole parameter space, a drastic reduction of $D_{KY}$ compared to a system with constant delay (black solid line, see text) is observed. Laminar chaos is found for $\lambda[f]+\lambda[R]<0$, i.e., whenever $\lambda[R]$ is below the dotted line in the center panels.
@@
In the following, we consider Eq. (1) with a quadratic nonlinearity f(z)=3.8z(1z)f(z)=3.8\,z(1-z) and the random and chaotically time-varying delays illustrated in Fig. 1 and Fig. 2(b), respectively. For the numerical analysis we choose $\Theta=200$ such that the small time scale $(1/\Theta)$ of the system is much smaller than the delay $\tau(t)$ and also much smaller than the time scale of the temporal structures of the delay variation. The results are presented in Figs. 3 and 5 for random and chaotically time-varying delays, respectively. For highlighting the differences between these and regular delay variations, analogous results for a periodic (sinusoidal) delay variation are shown in Fig. 4. Since laminar chaos and its generalization can only be observed in systems with dissipative delay, our first step is the classification of the time-varying delays in the delay parameter space using the Lyapunov exponent $\lambda[R]$ of the access map. For general one-dimensional maps, the Lyapunov exponent is defined by [107]
@@
To highlight the differences that come with random and chaotically time-varying delay, we first consider a generic example of a regular delay variation: The sinusoidal delay given by Eq. (2). Its parameter space shown in the top panel of Fig. 4 is characterized by two dynamical regimes of the access map and its corresponding circle map, Eq. (16) with $L=1$. Inside the so-called Arnold tongues [114, *arnold_small_1961_erratum], which are the dark tongue-like structures where $\lambda[R]$ is negative, the circle map shows stable periodic dynamics and the resulting delays are classified as dissipative delays. The complement of the Arnold tongues, which for fixed $A$ is a Cantor set with nonzero measure, i.e., a fat fractal [107], leads to marginally stable quasiperiodic dynamics of the circle map, $\lambda[R]=0$ and thus the resulting delay are classified as conservative delays. As illustrated in the center panel, $\lambda[R]$ as a function of $\tau_{0}$ for fixed $A=0.9$ also reflects this fractal structure showing an irregular behavior while it is periodic in $\tau_{0}$, where the period equals the delay period. For the delay system this means that the dynamics changes irregularly between high-dimensional turbulent chaos, and low-dimensional (generalized) laminar chaos if the delay parameters are varied. This has drastic consequences for the Kaplan-Yorke dimension $D_{KY}$ as shown in the bottom panel for fixed $A=0.9$. If the mean delay $\tau_{0}$ is varied, one observes huge variations of $D_{KY}$ over several orders of magnitudes, where the local maxima are roughly proportional to $\tau_{0}$. The overall variation of $D_{KY}$ increases with the parameter $\Theta$ since the local maxima are also proportional to $\Theta$ as they correspond to conservative delays leading to turbulent chaos, where the system is equivalent to a constant delay system so that the theory in [74] applies. In contrast, the local minima hardly vary with $\Theta$ given that $\Theta$ is large enough since the system is then well described by the low-dimensional dynamics of the limit map, Eq. (12). Such huge variations of $D_{KY}$ in a fractal manner are also observed for quasiperiodic delays [54].
@@
The situation drastically changes, when it comes to a random delay variation as shown in Fig. 3. While for small values of the mean delay $\tau_{0}$ we still find fractal structures in the parameter space similar to the ones observed for periodic delays, the fractal structures disappear for larger $\tau_{0}$. Moreover, the Lyapunov exponent of the access map seems strictly negative for all $A>0$ if $\tau_{0}$ exceeds a certain threshold, which indicates that almost all parameters lead to dissipative delays. This is also confirmed in the center panels, where $\lambda[R]$ is shown as a function of $\tau_{0}$ for fixed $A=0.9$. According to the theory on laminar chaos in Sec. III, the delay system defined by Eq. (1) shows low-dimensional (generalized) laminar chaos for almost all delay parameters, when $\tau_{0}$ exceeds a certain threshold. In Sec. IV.2 we find that this threshold is of the order of the correlation length of the delay variation. The low-dimensionality of the dynamics of the delay system is confirmed by the bottom panels, where we find that the Kaplan-Yorke dimension $D_{KY}$ for random delays (blue line) with large enough $\tau_{0}$ is of orders of magnitudes smaller than the dimension observed for the comparison system with a constant delay (black line). This was already conjectured based on results for a quasiperiodic delay with a large number of frequencies [54] and is now confirmed.
@@ IV.2 Theory
In order to explain the strictly negative value of the Lyapunov exponent of the access map that is observed in Fig. 3 and Fig. 5 for larger values of $\tau_{0}$, we follow the argumentation on disordered circle maps in the limit $L\to\infty$ found in [94]. There the finite-time Lyapunov exponent $\lambda_{K_{L}}[R_{\omega}]=\frac{1}{K_{L}}\sum_{k=0}^{K_{L}-1}\ln(1-\dot{\tau}_{\omega}(\theta_{k}))$ of an orbit $\mathcal{O}_{L}=\{\theta_{k}\}_{k}=0,1,\dots,K_{L}-1$ that passes the phase space of the circle map once is considered. Assuming that the minimal delay $\tau_{\text{min}}=\min_{\omega,t}\tau_{\omega}(t)$ is larger than the correlation length $l$ of the delay $\tau_{\omega}(t)$, $\tau_{\text{min}}>l$, one finds for the delays defined in Sec. II that the increments $\tau_{\omega}(\theta_{k})$ and $\tau_{\omega}(\theta_{k+1})$ are nearly independent. More precisely, if the tails of the the Gaussian $G_{\varsigma}(t)$ in Eq. (7) are neglected and therefore the Gaussian is replaced by a distribution with finite support [l,l][-l,l] the increments are strictly independent. In this case, the circle map with quenched disorder defined by Eq. (20) is stochastically equivalent to the circle map with annealed disorder, where the disorder changes with each iteration $\omega=\omega_{k}$. In the literature, such maps are known as “random dynamical system on the circle” [119, 120] (see [94] for more references) and it is known that their Lyapunov exponent is strictly negative if the realizations of the circle maps obtained from the disorder realizations $\omega_{k}$, almost surely, do not preserve the same measure [119, 120], which is a very general assumption. This means that $\lambda_{K_{L}}[R_{\omega}]$ converges to a unique, strictly negative value $\lambda[R]$ in the limit $L\to\infty$, which is associated with $K_{L}\to\infty$ as the length $K_{L}$ of the orbit $\mathcal{O}_{L}$ grows with the system size, $K_{L}\sim L/T$, where $T$ is the limit of the average roundtrip time, Eq. (22), for $L\to\infty$. For finite $L$, the authors of [94] further consider the distribution of the Lyapunov exponent of the disordered circle map, Eq. (20), with respect to the disorder. Using the approximate independence of subsequent increments $\tau_{\omega}(\theta_{k})$ and the central limit theorem, they find that the probability of generating a disorder realization with a Lyapunov exponent larger than $\lambda_{\text{cutoff}}$ with $\lambda_{\text{cutoff}}>\lambda[R]$ vanishes for increasing $L$ with
@@ V.1 Test for laminar chaos
Figure 7: Return maps of the levels of the laminar phases. The level $z_{n+p}$ of the $(n+p)$th laminar phase is plotted over the level $z_{n}$ of the $n$th laminar phase for different pp. The $z_{n}$ were extracted from a laminar chaotic time series of Eq. (1) with the chaotically time varying delay shown in Fig. 2(b), where we set $\tau_{0}=3$ and $A=0.9$. While non of the return maps resembles the nonlinearity of Eq. (1) (dashed lines), remnants of the nonlinearity are visible for $1\leq p\leq 4$, which indicates that $p$ is time-dependent and varies between these values for the given parameters.
Figure 8: Analysis of the nonlinear correlations in the return maps in Fig. 7 of a laminar chaotic time series. A value of (1Q){(1-Q)} and the MIC close to one would indicate strong nonlinear correlations as one would expect for at least one value of $p$ in the case of laminar chaos in systems with (quasi-)periodic delay. For random and chaotically time-varying delays, much smaller values are observed for all $p$, which indicates only weak correlations. This supports our conjecture that $p$ is time-dependent in this case.
@@
Figure 11: Return map of the levels $z_{n}$ of the laminar phases, where the sequence of the $p_{n}$ was extracted from the time series without knowing the system but using the clusters identified in Appendix A, Fig. 16. The nonlinearity (dashed line) is well resembled, but we find more outliers compared to Fig. 10, where the delay function was used to determine the sequence of the $p_{n}$. This is also reflected by slightly lower values of the quantities $(1-Q)\approx 0.79$ and $\text{MIC}\approx 0.77$, which nevertheless indicate strong nonlinear correlations as expected for laminar chaos.
In the following we naively apply the original toolbox to a time series of a system with chaotically time-varying delay, where we choose the same parameters as in Fig. 5 and Fig. 6(d)  [^7] The time series has a total length of $10^{4}$ time units, where we skip the first $10^{3}$ time units to let the transients relax. . We first identify the laminar phases and extract their levels $z_{n}$  [^8] To identify the laminar phases, we first detect the bursts between the laminar phases, where a burst is assumed when the absolute valued derivative $|\dot{z}(t)|$ of the time series exceeds the threshold $0.1$ and the duration between the end of the preceding burst and the current burst candidate is larger than $0.05$. So a laminar phase is detected if $|\dot{z}(t)|$ is close to zero for a duration that is at least one magnitude larger than the characteristic time scale $1/\Theta$ with $\Theta=200$. The burst position is then given by the onset of the burst, where $|\dot{z}(t)|$ starts to exceed the threshold. The level $z_{n}$ of the $n$th laminar phase is defined as the value of $z(t)$ in the middle between two subsequent burst positions.. Then the return map of the laminar phases is plotted, which should reveal the nonlinearity $f$ of the delay system if laminar chaos is present. However, this concept is only partially applicable for chaotically time-varying delay as shown in Fig. 7. The return maps clearly show remnants of the nonlinearity $f$ but with a noisy background and with remnants of higher iterates $f^{l}$. While for periodic and quasiperiodic delay the first iteration of the map $z^{\prime}=f(z)$ appears for a unique value $p=p_{\text{true}}$, it is found in Fig. 7 for $p=1,2,3,4$ (dashed line). This indicates that the number $p_{\text{true}}$ of laminar phases per solution segment is in general not constant for random and chaotically time-varying delay. As done in [95], we now quantify the nonlinear correlations of these return maps, where strong nonlinear correlations are an indicator for the presence of laminar chaos generated from systems with periodic or quasiperiodic delay. For that we compute the maximal information coefficient (MIC) introduced in [129] using the Python package minepy [130] and the quantity (1Q)(1-Q), which relates local fluctuations of the return map to global fluctuations of the levels $z_{n}$  [^9] Following the idea of the CANOVA method [143], we relate local fluctuations of the return maps of the $z_{n}$ to their global fluctuations. Therefore we sort the sequence of points $(z_{n},z_{n+p})$ by the first component leading to the sequence of points $(v_{m},w_{m})$. The quantity $Q$ is then defined by $Q=\frac{\text{Var}(w_{m+1}-w_{m})}{2\,\text{Var}(w_{m})}$, where $\text{Var}(w_{m})$ is the sample variance of all values $w_{m}$. More details can be found in [95]. . The results for our system with chaotic delay are shown in Fig. 8, where values close to 0 or $1$ belong to weak or strong correlations, respectively. Since both quantities are rather closer to 0 than to $1$, the classical test for laminar chaos fails. The reason for that is illustrated in Fig. 9, where part of the analyzed time series is shown together with the numerically detected boundaries of the laminar phases and with the boundaries of the state intervals $\mathcal{I}_{k}=(t_{k-1},t_{k}]=(t_{k}-\tau(t_{k}),t_{k}]$, which are the domains of the solution segments $z_{k}(t)$ representing the memory of the delay system at time $t=t_{k}$. Highlighted by the different shadings, the number $p=p_{k}$ of laminar phases per solution segment varies, so that the original test for laminar chaos fails as $p$ is assumed constant. If the delay function $\tau(t)$ is known, it is straightforward to obtain the sequence $p_{n}$ such that the points $(z_{n},z_{n+p_{n}})$ resemble the nonlinearity $f$. Therefore we assign to the $n$th laminar phase a time $\tilde{t}_{n}$ equal to the numerically detected center of the phase. Using the monotonicity of the access map, we know that the $n$th laminar phase is uniquely mapped to the $(n+p_{n})$th laminar phase. It follows that the correct $p_{n}$ with $z_{n+p_{n}}=f(z_{n})$ is found if $R(\tilde{t}_{n+p_{n}})$ is inside the domain of the $n$th laminar phase. Applying this approach to our exemplary time series gives us the associated sequence $p_{n}$. However, for some $n$ no or no unique $p_{n}$ can be detected. This happens when the plateau detection algorithm fails to detect a burst between two laminar phases, when spurious laminar phases are detected, or naturally when the number of laminar phases per solution segment changes, $p_{n+1}\neq p_{n}$ (see transitions between the regions with different color in Fig. 9). Plotting the points $(z_{n},z_{n+p_{n}})$ while omitting points with missing $p_{n}$ results in the return map shown in Fig. 10, clearly reproducing the nonlinearity except for some outliers and showing strong nonlinear correlations as expected for laminar chaos.
Figure 12: Mechanism behind laminar chaos leads to characteristic structures in $d$-dimensional return maps of the durations of the laminar phases with (a) $d=2$, (b) $d=3$. Tuples of durations $(\delta_{n},\delta_{n+1},\dots)$ of laminar phases with the same number $p_{n}=p_{n+1}=\dots=p$ of laminar phases inside the solution segments belong to the same point cluster. Depending on $p$ these clusters scatter around (p1)(p-1)-dimensional affine subspaces, which can be used to reconstruct the sequence $p_{n}$ from time series without knowing the system (see text). In (a), $p=1$, $p=2$, $p=3$, $p>3$ are represented by the blue squares, orange triangles, green diamonds, and light gray dots, respectively. If no unique $p$ is numerically found, dark grey crosses are used. In (b), the color coding is same and the (p1)(p-1) subspaces defining the shape of the cluster are illustrated for $p=1,2,3$ by the black dot, line, and plane, respectively.